If A and B are squares matrices such that A^2006=0 and AB=A+B,then find det(B).

Grade:12

1 Answers

mycroft holmes
7 years ago
We have det A = 0. Also from A+B = AB we have B = AB-A = AB – AI = A(B-I)
 
so det B = det A det(B-I) = 0